OpenStudy (anonymous):

If 2.078 g of Smarties® requires 2.11 mL of 0.250 M Na 2CO3 to reach the endpoint, how many moles of acid is there in each gram of the candy?

OpenStudy (anonymous):

greetings do you understand any of this s** sorry but do you? I have an equation with solution if you wanna check it over...see if it's right?

OpenStudy (abmon98):

please do :)

OpenStudy (anonymous):

(2 x .25 x 2.11)/3 = 3.51 moles per gram

OpenStudy (abmon98):

Number of Moles=Concentration(mol/dm^3)*Volume(dm^3) Moles of NaOH=2.11*0.250/1000=0.0005275 Is there any balanced equation provided that would be useful

OpenStudy (anonymous):

umm... hold on... (randomly flipping through lab papers) 2 HOC(CO2H)(CH2C)2H)2 + 3 NA2CO3 -> HOC(CO2^-Na^+)2 + 3 CO2 + 3H2O

OpenStudy (abmon98):

i made a mistake Na2CO3** look at the mole ratio of 2 HOC(CO2H)(CH2C)2H)2 to 3 NA2CO3 for every 2 moles of smarties we need 3 moles of sodium carbonate. 2:3 x:0.0005275 x=0.0005275*2/3 Number of moles=Mass(g)/Molar Mass(g/mol) Mass is given and number of moles are given workout to get the molar mass.

OpenStudy (anonymous):

umm...ok so what I got from your "smarty pants science speak" was my equation was wrong, and I need to do more work *drama queen sigh* very well...back to the trenches I go

OpenStudy (abmon98):

lol,i am sorry :/

OpenStudy (anonymous):

Wait...no...hold on. I didn't give you their equation for calculating molar acidity or whatever they call it un momento :)

OpenStudy (anonymous):

moles of acid = a x M x Vb a = Reaction coefficient of the acid Mb = Molarity of the basic solution Vb = Volume of base used in titration converted to Liters b = Reaction coefficient of the base

OpenStudy (anonymous):

is that at all helpful? sorry bout that

OpenStudy (abmon98):

i do not about this equation but looks like all your givens are provided so whats left is to plug it.

OpenStudy (abmon98):

but i think the number moles of acid is based on the 2.078 grams of smarties not 1 gram

OpenStudy (anonymous):

yeah...wait... hold on

OpenStudy (anonymous):

I don't..I don't...just look! sorry I suck at chemistry, which makes no sense, as I love math. Chemistry just irritates the sh** out of me.

OpenStudy (abmon98):

Calm down :) In the link you have posted it says that smarties (citric acid) 2 HOC(CO2H)(CH2CO2H)2 + 3 Na2CO3--->HOC(CO2- Na+)(CH2CO2- Na+)2 3CO2 + 3H2O Acid is a proton Donor and base is a proton acceptor. Citric acid consists of (CH3COOH)3 Acid+Base(Carbonate)-->Carbon Dioxide +Water+Salt Citric Acid donates 3 H+ to form 3 moles of H2O as they will require 3OH- to accept the H+. HOC(CO2H)(CH2CO2H)2+ we have 3 structure of CH2CO2- they are attracted to 3Na+ as the reaction progresses. thats the reason of why 3 moles of Na2CO3 are required to react with 2 moles of citric acid. look at the balanced equation above. 2.11 mL of 0.250 M Na 2CO3 require 2.078 g of smarties to react. In order to find the number of moles of NaOH we use the equation Number of moles=Concentration(mol/L)*Volume(L) We change the Volume from ml to L by dividing by 1000 to get equal units with concentration. 2.11/1000*0.250=0.0005275 moles of NaOH From the balanced equation in order for the reaction to occur completely it requires 2 moles of citric acid(smarties) to react with 3 moles of NaOH what if 0.0005275 moles of Na2CO3 are used, we use mole ratio \[2:3\] \[x:0.0005275\] x represent the unknown moles of citric acid. cross multiplication and you can find x easily. \[x=0.0005275*2/3=0.0003516667 mole\] mole per gram simply means the molar mass Number of moles=Mass(g)/Molar Mass(g/mol) Mass is given in your question which is 2.078 g of smarties and now you,ve got the number of moles rearrange the equation to Molar Mass(g/mol)=Mass(g)/Number of Moles Molar Mass(g/mol)=2.078/0.0003516667=5909.00473